[figure: 9.7 cm]

SPIN-WAVES IN MAGNETS WITH IMPURITIES: LANDAU-LIFSHITZ EQUATION APPROACH

E.G. Mishchenko

Lorentz-Instituut, Leiden University, P.O. Box 9506, 2300 RA Leiden, The Netherlands and Landau Institute for Theoretical Physics, Kosygin 2, 117334 Moscow, Russia

Using the Landau-Lifshitz equation the interaction of long-wavelength excitations with magnetic defects in ferro- and antiferromagnets is studied. The defects are supposed to produce random inhomogeneous magnetic field and may have different geometry, e.g. be point, line or plane imperfections. The static random magnetic structure caused by this field is determined together with equations for the spin waves. The Dyson equation for a magnon self-energy is derived and its solution for different scales of inhomogeneities is analysed. The magnon frequency shift and damping are found as well as the lineshape of ferromagnetic resonance. It is shown that the width of the uniform resonance has an essentially nonlinear dependence on the impurity concentration, the effect missed by the usual first Born approximation theory. The impurities in an antiferromagnet are also considered using the same approach. This work was supported by the NWO/FOM Dutch Science Foundation and by the Russian Foundation for Basic Research (Grant No. 97-02-16044).

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : E.G. Mishchenko

Address for correspondence : Lorentz-Instituut, Leiden University, P.O. Box 9506, 2300 RA Leiden, The Netherlands

Email address : mishch@lorentz.leidenuniv.nl